OpenStudy (anonymous):

Find the greatest common factor of the pair of numbers. 333 and 441

OpenStudy (tkhunny):

Prime Factorization!!!!! 333 = 3 * 111 = 3 * 3 * 37 441 = 3 * 147 = ??

OpenStudy (anonymous):

9 333=3*3*37 441=3*3*7*7